Diffusion Morning Report 05/20/04

ted berg tberg at snf.stanford.edu
Thu May 20 06:57:40 PDT 2004


Good Morning All,
    Looking pretty good this morning.  LTO is back up after tube change. 
Nitride is under observation. Poly is getting close to pull and
clean, we will take it down on Tuesday.  Epi Has water issue again- 
looking into it. TEOS 2 still has generator issues. Everything else
looks OK.

Diffusion 	Date: 	5/20/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	MFC is installed testing-trap installed 	 
Tylan 8-SiGe 	Up 	  	6.6/30
Tylan 9-Poly 	Up 	  	29.17/30
Tylan 10-Nitride 	Up 	Particles? 	8.3/30
Tylan 11-BPSG-LTO 	Up 	test good 	.6/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	11.6/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 520 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040520/76f12c79/attachment.html>


More information about the diffusion mailing list